Given a transformation  defined as. Find  and , range of                                                                


1
Expert's answer
2021-09-02T10:15:12-0400

"Let L: R^{3} \rightarrow R^{2} \ be\ defined \ by:\\
 
L\left(\begin{array}{l}
x_{1} \\
x_{2} \\
x_{3}
\end{array}\right)=\left(\begin{array}{l}
x_{1}+x_{2} \\
x_{2}+x_{3}
\end{array}\right)\\
 
\therefore{ker} L= 
 
\begin{aligned} &\operatorname{ker} L= \\
&\left\{\left(\begin{array}{l}
x_{1} \\
x_{2} \\
x_{3}
\end{array}\right) \mid\left(\begin{array}{l}
x_{1}+x_{2} \\
x_{2}+x_{3}
\end{array}\right)=\left(\begin{array}{l}
0 \\
0
\end{array}\right)\right\}
\end{aligned}
 \\
The \ homogeneous \ system \ coefficient \ matrix:
 
\left(\begin{array}{lll}
1 & 1 & 0 \\
0 & 1 & 1
\end{array}\right) \text { has r.r.e.f. }\left(\begin{array}{ccc}
1 & 0 & -1 \\
0 & 1 & 1
\end{array}\right)
 
 x_{3} \ is \ arbitrary, \\ x_{1}=x_{3}, x_{2}=-x_{3} \\
 \operatorname{ker} L=\operatorname{span}\{\underbrace{(1,-1,1)}\} \\ 
 \\\\\\ \quad basis \ for \ ker L. \\
\\\operatorname{ker} L \neq\{(0,0,0)\}\\
 
\begin{aligned}
&\text { range } L=\left\{\left(\begin{array}{l}
x_{1}+x_{2} \\
x_{2}+x_{3}
\end{array}\right) \mid \text { for all } x_{1}, x_{2}, x_{3}\right\} \\
&=\left\{x_{1}\left(\begin{array}{l}
1 \\
0
\end{array}\right)+x_{2}\left(\begin{array}{l}
1 \\
1
\end{array}\right)+x_{3}\left(\begin{array}{l}
0 \\
1
\end{array}\right) \mid\right. \text { for } \\
&\text { all } \left.x_{1}, x_{2}, x_{3}\right\}
\end{aligned}\\
 
Find \ a \ basis \ for \ range \ L= span\{(1,0),(1,1),(0,1)\}
 \\
 
\begin{aligned}
&\left(\begin{array}{lll}
1 & 1 & 0 \\
0 & 1 & 1
\end{array}\right) \text { has r.r.e.f. }\left(\begin{array}{lll}
1 & 0 & -1 \\
0 & 1 & 1
\end{array}\right) \\
&\Rightarrow \text { range } L=\text { span }\{\underbrace{(1,0),(1,1)}\}
\end{aligned}\\
 
range L=R^{2} \\\Rightarrow L \ is \ onto."